Jan's age is 3 years less than twice Tritt's age. The sum of their age is 30. Find this ages.
Answer by macston(5194)

.
T=Tritt's age; J=Jan's age=2T-3yrs
.
T+J=30yrs
T+(2T-3yrs)=30yrs
3T=33yrs
T=11yrs
ANSWER 1: Tritt is 11 years old.
J=2T-3yrs=2(11yrs)-3yrs=22yrs-3yrs=19yrs
ANSWER 2: Jan is 19 years old.